Lane invested $22,000, part at 8% and part at 7%. If the total interest at the end of the year is $1,710, how much did she invest at 8%?
Answer by checkley75(3666)

.08X+.07(22,000-X)=1,710
.08X+1,540-.07X=1,710
.01X=1,710-1,540
.01X=170
X=170/.01
X=17,000 AMOUNT INVESTED @ 8%.